OpenStudy (anonymous):

Find the exact value of tan 22.5 by using the expression for tan2x. Medal will be given for best ans

OpenStudy (anonymous):

11.25

OpenStudy (anonymous):

The textbook answer is apparently: \[\sqrt{2} -1\]

OpenStudy (sidsiddhartha):

use tan2a=2tana/(1-tan^2a) so tan45=2tan22.5/(1-tan^22.5) tan45=1 so tan^22.5+2*tan22.5-1=0 solve it

OpenStudy (sidsiddhartha):

sorry made a mistake tan45=2tan22.5/[1-tan^2(22.5)]

OpenStudy (anonymous):

tan 45 = 2tan22.5 / 1-tan^2 22.5 1 = 2tan22.5 / 1- tan^2 22.5 by cross multiplication u get: 1- tan^2 22.5 = 2tan 22.5 u get a quadratic and solving it you'll get 2 ans : -1+sqrt2 and -1-sqrt2 but only fist one is right cause 22.5 is in 1st quadrant and hence tan22.5 should be positive.So the second one will be wrong ans

OpenStudy (sidsiddhartha):

so it will look like so ur ans is =sqrt(2)-1